OpenStudy (anonymous):

help i ned help answers now for medals 5 questions

OpenStudy (anonymous):

Translate the sentence into an equation. Six less than a number is 14. A. 6 – n = 14 B. 6n = 14 C. n – 6 = 14 D. n = 14 – 6

OpenStudy (anonymous):

C :)

OpenStudy (anonymous):

4 more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Translate the sentence into an equation. Six less than the quotient of a number and 5 is 10. A. B. C. 5n – 6 =10 D. (n + 5) – 6 = 10 the firs atachement is b the secont is a

OpenStudy (anonymous):

Sorry It's my internet is being slow 1 min.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

I seriously hate my computer ugh. try this website www.wolframalpha.com it usually shows everything you need to know. The second one i'm not sure on. what are the other questions?

OpenStudy (anonymous):

Question Navigator -------------------------------------------------------------------------------- 3. Translate the sentence into an equation. Four times the difference of a number and 5 is 24. A. 4 • (c – 5) = 24 B. 4c – 5 = 24 C. c – 4 • 5 = 24 D. 5 • (c – 4) = 24

OpenStudy (anonymous):

do you want to try to learn how to do it? Or just the answer, I want you to actually learn haha

OpenStudy (anonymous):

ok i dc

OpenStudy (anonymous):

ok haha. Do you understand how i got the first answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok Well it says Six less than a number is 14. Since we do not know what the number is, we use n as the variable. We have to subtract 6 from n. so that is n – 6 = 14. =14, because we know that's what we have to end up with.

OpenStudy (anonymous):

so 6 +14=n

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay so wat is this

OpenStudy (anonymous):

well you try to figure out the third one. I want to make sure you get it

OpenStudy (anonymous):

4 times c +5 = 24 whitch means u have to divide 24 by 4 whitch is 6 and then subtract 5 from 6 whitch is 1 c = 1

OpenStudy (anonymous):

Four times the difference of a number and 5 is 24. so I think you're right.

OpenStudy (anonymous):

yayt so was awnser??

OpenStudy (anonymous):

I think it's A. Because difference is subtraction. Sorry I forgot to point that out.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Choose the equation that models the problem. Zach's baseball team won 32 games last season. They played a total of 45 games. How many games did they lose overall? Let l represent the number of games that Zach's team lost. A. l – 32 = 45 B. l + 32 = 45 C. l + 45 = 32 D. 32l = 45

OpenStudy (anonymous):

Zach's baseball team won 32 games last season. They played a total of 45 games. How many games did they lose overall? Let l represent the number of games that Zach's team lost. Ok so it's B. We have to add the amount they lost and won to get the total, l + 32 = 45

OpenStudy (anonymous):

Solve the word problem. Olivia has 21 coins in her drawer. Nine of the coins are nickels, and the rest are pennies. What is the total value of the coins in her drawer? A. 21¢ B. 30¢ C. 45¢ D. 57¢

OpenStudy (anonymous):

We have to do 9*5 first to get how much the nickels are worth which is 45. Now we can subtract 9 from 21. so we have 12 coins that are pennies. So 45+12=57. So the answer is D. 57c

OpenStudy (anonymous):

thnx i got 80 %

OpenStudy (anonymous):

Awesome. Which one was wrong?